### v1

# Communication-Efficient Decentralized Optimization via Double-Communication Symmetric ADMM

Abstract
This paper focuses on decentralized composite optimization over networks without a central coordinator. We propose a novel decentralized Symmetric ADMM algorithm that incorporates multiple communication rounds within each iteration, derived from a new constraint formulation that enables information exchange beyond immediate neighbors. While increasing per-iteration communication, our approach significantly reduces the total number of iterations and overall communication cost. We further design optimal communication rules that minimize the number of rounds and variables transmitted per iteration. The proposed algorithms are shown to achieve linear convergence under standard assumptions. Extensive experiments on regression and classification tasks validate the theoretical results and demonstrate superior performance compared to existing decentralized optimization methods. To our knowledge, this is the first decentralized optimization framework that achieves a net reduction in total communication by leveraging fixed multi-round communication within each iteration.
1 Introduction
The increasing size and complexity of modern machine learning models, combined with the explosive growth of data from sources such as mobile devices, sensors, and edge computing platforms, has driven the demand for scalable and privacy-preserving optimization techniques. Among these, decentralized optimization has emerged as a powerful approach, particularly when centralized computation is impractical due to concerns of scalability, robustness, data privacy, communicational infeasibility and network connectivity.
Unlike centralized distributed optimization, which still depends on a central server to coordinate updates, decentralized optimization involves multiple agents collaboratively solving a global problem by performing local computations and exchanging information only with their neighbors. These agents operate over a connected network—typically modeled as a graph—without the need for a central coordinator. This architecture makes decentralized optimization especially attractive for applications in multiple fields like sensor networks and large-scale machine learning.
In decentralized optimization, a central challenge lies in reducing the time cost of both local computation and inter-node communication. While existing algorithms differ in their local update strategies, most follow a common structural pattern: each iteration is followed by a single round of communication. This convention has seldom been challenged, primarily due to the concern that introducing multiple communication rounds per iteration would increase the overall communicational cost. Prior attempts to incorporate fixed multiple communication rounds, such as those in (Multi1, ; DGDT, ; ProxMudag, ; Multi2, ), achieve this by multi-consensus—repeatedly mixing local variables through communication. However, these methods have not demonstrated practical reductions in the total number of communication rounds. As a result, the potential for achieving a net communication reduction through non-adaptive multi-communication algorithms remains largely unexplored.
Although multi-consensus schemes involve more frequent averaging steps, they do not necessarily reduce the overall communication cost. A potential reason is that these repeated consensus/communication steps primarily accelerate agreement among local variables, but offer limited improvement to the quality of each iteration. This observation motivates the need for a more principled and dedicated framework for multi-round communication, rather than simply applying multi-consensus in iteration.
In this paper, we investigate the potential of integrating multiple rounds of communication within a single iteration motivated by this observation. Rather than directly applying multiple mixing steps, we develop our algorithms by introducing linear constraints tailored for ADMM, which naturally embed multi-round communication into each iteration. To further enhance performance, we adopt a Symmetric ADMM (SADMM, ) framework to accelerate convergence. Although this design increases the per-iteration communication cost, it enables a significantly faster convergence, leading to a substantial reduction in the total number of iterations, computations, and overall communication required for convergence.
Our contributions are summarized as follows:
  • We propose DS-ADMM, a novel Symmetric ADMM-based decentralized composite optimization framework that incorporates multiple communication rounds into each iteration, leading to more efficient decentralized training.
  • We derive optimal communication rules in the proposed algorithms which successfully minimize the communication rounds and the amount of information transmitted per iteration.
  • We provide rigorous theoretical guarantees for the proposed method, including convergence and convergence rate analysis under standard and strong convexity assumptions.
  • We conduct extensive numerical experiments that validate our theoretical results and demonstrate the superior performance of our method compared to state-of-the-art algorithms in decentralized composite optimization by reducing both computational and communicational cost.
To the best of our knowledge, this is the first work to reduce the total communication cost by enabling fixed multiple communication rounds within a single iteration. Our results open a promising direction for decentralized optimization by revealing a new trade-off between the number of per-iteration communication rounds and overall convergence speed.
1.1 Related Work
Decentralized Optimization.
Decentralized optimization has gained significant attention in large-scale machine learning, particularly in scenarios where data is distributed across multiple agents or devices without a central coordinator.
A common strategy in these methods is the use of stochastic or deterministic mixing matrices to perform local averaging of variables across neighboring nodes, facilitating global consensus through communication. Early methods such as Decentralized Gradient Descent (DGD) (DGD1, ; DGD2, ; DGDConv, ) directly extend classical gradient-based algorithms to networked environments, laying the groundwork for later developments. However, DGD suffers from slow convergence and sensitivity to step size. To address these shortcomings, a line of gradient-tracking based algorithms has been developed, including EXTRA and PG-EXTRA (EXTRA, ; PG-EXTRA, ), NIDS (NIDS, ), SONATA (SONATA, ), and (GT1, ), which incorporate correction terms to estimate the average gradient across the network.
Several of these algorithms, including (PG-EXTRA, ; NIDS, ; SONATA, ) are designed for decentralized composite optimization problems with smooth-nonsmooth structure by embedding proximal gradient steps. And the unifying analysis in (unify, ) shows that a broad family of methods—including (PG-EXTRA, ; NIDS, ; GT1, )—can be understood through a single theoretical framework and established linear convergence under strong convexity assumptions.
In addition to improving per-iteration convergence rates, many recent methods incorporate acceleration techniques such as Nesterov acceleration, leading to further improvements in both computation and communication complexity including (Multi1, ; Multi2, ; DNGD, ; ProxMudag, ).
In parallel, a different family of methods focuses on dual-based formulations. These include (Dual1, ; Dual2, ; Dual3, ) and decentralized adaptations of the Alternating Direction Method of Multipliers (ADMM) (DADMM1, ; DADMM, ; DADMM3, ; DADMM2, ) which reformulate the problem with consensus constraints and alternate between primal and dual updates. A notable development in this line is (DPADMM, ), which construct constraints based on mixing matrix into an ADMM framework for decentralized composite optimization.
ADMM and Symmetric ADMM
The Alternating Direction Method of Multipliers (ADMM) is a powerful algorithmic framework for solving linearly constrained convex optimization problems with separable objective structures, and it does so without requiring smoothness assumptions. This characteristic makes ADMM particularly well-suited for composite optimization tasks involving non-smooth loss functions and regularizers. We refer readers to (ADMMEckstein, ; Boyd, ; Glowinski2014, ) for comprehensive overviews of the classical ADMM framework, and to (PADMM, ; nonergo, ; ADMMconv1, ; ADMMconv, ) for detailed convergence analyses of ADMM and its various extensions.
To improve convergence speed and numerical performance, Symmetric ADMM (S-ADMM) and its generalizations have been proposed in recent years (SADMM, ; GSADMM, ). These methods modify the standard ADMM iteration by introducing a symmetric primal-dual update structure, typically involving an additional intermediate update of the dual variable. This symmetric design allows for more balanced update dynamics between the primal and dual variables and often leads to improved practical performance. Convergence analyses for S-ADMM and its extensions have been established in works such as (GSADMMConv, ; PRSM, ).
Multi-Communication in Decentralized Optimization
Incorporating multiple communication rounds per iteration has been explored in decentralized optimization for different purposes, using either fixed or adaptive strategies. Early work by (DGDT, ) showed that fixed multi-communication inherits DGD’s convergence issues and incur high communication costs, while adaptive strategies—where communication rounds increase periodically—can achieve exact convergence, albeit requiring tuning or prior knowledge.
Later methods (Multi1, ; ProxMudag, ; Multi2, ) adopted fixed multi-communication to attain optimal theoretical communication complexity. However, empirical results in (ProxMudag, ) indicate that fixed multi-round schemes are unable to reduce total communication. This limitation motivates algorithmic designs that embed multi-communication within the problem structure, rather than treating it as an external enhancement.
1.2 Notation
We denote $\mathbf{1}_{m}\in\mathbb{R}^{m}$ as the vector of all ones, and $I_{m}$ as the $m\times m$ identity matrix. The nullspace and range space of a matrix $A$ are denoted by $\mathrm{null}(A)$ and $\mathrm{span}(A)$, respectively. For any symmetric, positive definite matrix $M$ and any vector $x$ of compatible dimension, we define the weighted norm as $\|x\|_{M}:=\sqrt{\langle Mx,x\rangle}$. The distance from a point $x$ to a set $S$ with respect to $M$ is defined as $\mathrm{dist}_{M}(x,S):=\inf\left\{\left\|x-s\right\|_{M}\,\middle|\,s\in S\right\},$ and we omit the subscript $M$ when $M=I$.
The proximal operator of a convex function $g(\cdot)$ with parameter $\lambda>0$ is defined as
$\text{Prox}_{\lambda g}(v)=\underset{x\in\mathbb{R}^{m}}{\arg\min}\left(g(x)+\frac{1}{2\lambda}\|x-v\|^{2}\right).$ (1)
We use it for the proximal operator of many common functions have explicit forms.
2 Preliminaries
2.1 Problem Setup
In this paper, we consider a network of $n$ agents collaboratively solving a decentralized composite optimization problem of the form:
$\min_{x\in\mathbb{R}^{d}}F(x)=\sum_{i=1}^{n}\left[f_{i}(x)+g_{i}(x)\right],$ (2)
where $f_{i}$ is a convex loss function and $g_{i}$ is a convex local regularizer both privately held by agent $i$. In cases involving a global regularizer $g$, we decompose it across the agents via $g_{i}=\lambda_{i}g$, ensuring that $\sum_{i=1}^{n}\lambda_{i}=1$. This framework models many decentralized machine learning tasks. Typical examples of loss functions include least squares, quantile loss, Huber loss, and hinge loss, while examples for regularization include the $\ell_{1}$-norm, $\ell_{2}$-norm, and elastic net.
2.2 Graph Topology
The communication structure among agents is modeled by an undirected and connected graph $G=(V,E)$, where $V=\{1,2,\dots,n\}$ is the set of agents, and an edge $(i,j)\in E$ indicates a direct communication link between agents $i$ and $j$.
The graph structure is encoded by a mixing matrix $W\in\mathbb{R}^{n\times n}$, where $W_{ij}\in[0,1]$ denotes the communication weight between agents $i$ and $j$. The mixing matrix satisfies the following assumptions:
Assumption 1.
(1) $W$ is symmetric; (2) $W$ is doubly stochastic as $W\mathbf{1}=\mathbf{1}$, where $\mathbf{1}$ denotes the all-ones vector; (3) $W_{ij}>0,i\neq j$ if and only if $(i,j)\in E$ and $W_{ii}>0$ for all $i\in V$.
The assumption of the graph and corresponding mixing matrices leads to the following properties which can be derived from the Perron-Frobenius Theorem:
Proposition 1.
(1) The eigenvalues of the mixing matrix satisfy $1=\lambda_{1}(W)>\lambda_{2}(W)\geq\dots\geq\lambda_{n}(W)>-1$, and the spectral gap $\rho=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}>0$; (2) $\mathrm{null}(I_{n}-W)=\mathrm{span}\{\mathbf{1}_{m}\}$ .
See Appendix A for an example of mixing matrix based on Metropolis-Hastings weight (Weight, ).
3 Method
This section is of the design of our communication-efficient decentralized algorithm. We begin by reformulating the consensus constraint to accommodate multi-round communication within each iteration. We then describe how proximal linearization ensures decentralized updates, outline the communication strategy that minimizes per-iteration transmission, and finally present the full decentralized version of our proposed algorithm.
3.1 Reformulating Consensus Constraints
Which is crucial in applying Symmetric ADMM in decentralized optimization is to enforce agreement among all agents’ local variables. However, unlike in Distributed ADMM (Boyd, ) where a global coordinator can explicitly manage consistency, the decentralized setting requires a different strategy to ensure consensus, particularly to form suitable linear constraints. The inspiration was from the null space property of the matrix $I_{n}-W$. Define $\tilde{W}=W\otimes I_{d}$, where $\otimes$ denotes the Kronecker product. Let $u=(u_{1}^{\top},\dots,u_{n}^{\top})^{\top}\in\mathbb{R}^{nd}$ be the stacked vector of local variables across the network. It is straight to verify the following proposition since a positive spectral gap exist:
Proposition 2.
The consensus condition $u_{1}=u_{2}=\dots=u_{n}$ is equivalent to $(I_{nd}-\tilde{W}^{\top}\tilde{W})u=0$.
To integrate this consensus structure into our algorithm, we introduce the auxiliary constraint variable $v=(v_{1}^{\top},\dots,v_{n}^{\top})^{\top}\in\mathbb{R}^{nd}$ and propose the following symmetric formulation of the consensus constraint: $u=\tilde{W}v$, $\tilde{W}u=v$. This system embeds consensus into a linear structure that is both symmetric and decentralized, making it ideal for Symmetric ADMM application.
Denote $f(u)=\sum_{i=1}^{n}f_{i}(u_{i})$ and $g(v)=\sum_{i=1}^{n}g_{i}(v_{i})$, using these constraints, problem (2) can be equivalently reformulated as the following form:
$\underset{u,v\in\mathbb{R}^{nd}}{\min}f(u)+g(v)\quad s.t.\quad Au-Bv=0$ (3)
[…]
Remark 1.
The ADMM algorithm proposed in (DPADMM, ) adopts a similar constraint formulation, with $A=((I_{nd}-\widetilde{W})^{1/2},I_{nd})^{\top}$ and $B=(0,I_{nd})^{\top}$. This design enables a single communication round per iteration by absorbing part of the Lagrange multiplier update. However, this asymmetry prevents the application of Symmetric ADMM, which requires a balanced primal-dual formulation.
3.2 Proximal Linearization
[…]
This choice ensures that the augmented Lagrangian terms become separable across agents, enabling fully decentralized computation. It also aligns with prior work in proximal ADMM (PADMM, ), now adapted for use in a decentralized setting. As a result, the global update form (15) can be derived and presented in Appendix B for brevity.
Given the global update rules and the symmetry of the mixing matrix $W$, we can decompose the updates into per-agent computations and derive a fully decentralized algorithm.
3.3 Communication Strategy
For convenience, we define local aggregations via neighbor communication as $\tilde{w}_{1i}^{(t)}=\sum_{j=1}^{n}W_{ji}w_{1j}^{(t)}$, $\tilde{w}_{2i}^{(t)}=\sum_{j=1}^{n}W_{ji}w_{2j}^{(t)}$, $\tilde{v}_{i}^{(t)}=\sum_{j=1}^{n}W_{ji}v_{j}^{(t)}$, $\tilde{u}_{i}^{(t)}=\sum_{j=1}^{n}W_{ji}u_{j}^{(t)}$.
The use of the quadratic form $\widetilde{W}^{\top}\widetilde{W}$ in the global update rule (15) indicates that each agent’s update involves 2-distance neighbor information. This observation leads to the following result:
Proposition 3.
For our proposed algorithm, a single iteration which contains updates to both primal variables and two updates to the Lagrange multiplier requires at least two rounds of communication to transmit the necessary information.
Given the presence of four independent $d$-dimensional variables per agent, it is essential to optimize communication scheduling, both in timing and in which variables are transmitted. This leads to two guiding principles: (1) restrict communication to two rounds per iteration, and (2) minimize transmitted data per round.
Each Symmetric ADMM iteration follows the sequence: $u$-update → first $\lambda$-update → $v$-update → second $\lambda$-update. The first update of $w_{1}$ depends on mixed $u$ information, requiring a communication step between the $u$-update and the first $w_{1}$-update. Likewise, a second communication is needed after the $v$-update to support the $w_{2}$-update. These two rounds are essential and cannot be collapsed.
To reduce communication overhead, we carefully select what is transmitted. In the first round, sending updated $u$ is sufficient for updating $w_{1}$, but not for updating $v$, which depends on more refined information. Rather than transmitting primal variables directly, we send dual variables that implicitly encode the needed content. Specifically, transmitting $a_{i}^{(t)}=w_{2i}^{(t+\frac{1}{2})}+\frac{1}{r}(w_{2i}^{(t+\frac{1}{2})}-w_{2i}^{(t)})$ allows correct computation of $v_{i}^{(t+1)}$. This positions the first communication round between the first $w_{2}$-update and the $w_{1}$-update. The second round similarly transmits $b_{i}^{(t+1)}=w_{1i}^{(t+1)}+\frac{1}{s}(w_{1i}^{(t+1)}-w_{1i}^{(t+\frac{1}{2})}).$ between the second $w_{1}$-update and the $w_{2}$-update.
Together, these insights determine the optimal placement and content of communication in each iteration, with two rounds of communication per iteration and two $d$-dimensional variables transmitted per round.
Crucially, our design partitions the variables into two interdependent groups: $(u,w_{2})$ and $(v,w_{1})$. Information from one group is not used directly in its own update but enables the update of the other group, creating a feedback structure. This interleaving induces a coupled communication-update mechanism where each block drives progress in the other.
Furthermore, this structure makes Symmetric ADMM not just suitable but essential: it provides accelerated convergence without increasing communication, and leads to a clean, symmetric algorithm. To our knowledge, this tightly coupled update-communication framework is novel in the decentralized optimization literature.
Remark 2.
Dividing the Lagrange multiplier $\lambda$ into two independent blocks $w_{1}$ and $w_{2}$ greatly enhances flexibility in communication design since the updates of two blocks are parallel to each other and realigning the order of them is allowed.
3.4 Algorithm
We denote our proposed algorithm as DS-ADMM, which encapsulates several key design features: it adopts the Symmetric ADMM framework for Decentralized optimization, incorporates a Double communication structure within each iteration, and exhibits a Symmetric structure in its use of two interleaved variable blocks.
We now present the algorithm in its fully decentralized form. For clarity, we reposition the update of the dual variable $w_{2}$—originally placed at the end of each iteration—to the beginning. The step-size parameters are set as $0<r\leq 1$ and $s=1$.
Algorithm 1 DS-ADMM
1:Initialize: $u_{i}^{(0)}=v_{i}^{(0)}=w_{1i}^{(0)}=w_{2i}^{(-\frac{1}{2})}=0$ for all $i\in\{1,\dots,n\}$ for all $i\in\{1,\dots,n\}$, mixing matrix $W\in\mathbb{R}^{n\times n}$.
2:repeat
[…]
4:  [Communication 1 ] Transmit $a_{i}^{(t+1)}=w_{2i}^{(t+\frac{1}{2})}+\frac{1}{r}(w_{2i}^{(t+\frac{1}{2})}-w_{2i}^{(t)})$ and $u_{i}^{(t+1)}$.
[…]
6:  [Communication 2] Transmit $v_{i}^{(t+1)}$ and $b_{i}^{(t+1)}=2w_{1i}^{(t+1)}-w_{1i}^{(t+\frac{1}{2})}$.
7:until convergence criterion is satisfied
4 Convergence Analysis
In this section, we analyze the convergence properties of the proposed decentralized algorithm. As it is a direct application of Symmetric ADMM with proximal terms, several convergence results follow from existing literature. We further establish linear convergence under specific conditions that are mild yet broadly applicable to machine learning problems.
To facilitate the analysis, we define the block matrix:
$H=\begin{pmatrix}Q&&\\ &Q+\frac{1}{r+1}\beta B^{\top}B&-\frac{r}{r+1}B^{\top}\\ &-\frac{r}{r+1}B&\frac{1}{\beta(r+1)}I\end{pmatrix}$, (4)
which is positive definite. We also define the concatenated variable $w=(u^{\top},v^{\top},\lambda^{\top})^{\top}\in\mathbb{R}^{4nd}$.
4.1 General Sublinear Convergence
Theorems 3.3 and 4.2 of (PRSM, ) imply that the proposed algorithm enjoys a general sublinear convergence rate $\mathcal{O}(1/t)$ without requiring strong assumptions on the objective functions.
Theorem 1.
Let $\{w^{(t)}\}$ be the sequence generated by DS-ADMM. Then $\{w^{(t)}\}$ converges to a solution point $w^{\infty}$, and the following non-ergodic sublinear rate holds:
$\|w^{(t)}-w^{(t+1)}\|^{2}\leq\frac{1}{\beta\tau(t+1)}\cdot\frac{1+r}{1-r}\left(\|w^{(1)}-w^{(0)}\|_{H}^{2}+\|v^{(1)}-v^{(0)}\|_{Q}^{2}\right).$ (5)
Remark 3.
This sublinear convergence result is independent of the underlying communication graph and mixing matrix. Thus, the algorithm is inherently robust to different network topologies in decentralized environments.
4.2 Linear Convergence under Metric Subregularity
Although various results on linear convergence of ADMM and its variants exist (e.g., (PADMM, ; ADMMconv1, ; ADMMconv, ; GSADMMConv, ; PRSM, )), none directly apply to our algorithm. Nevertheless, we adapt ideas from these works to establish linear convergence under a standard regularity condition known as metric subregularity.
Definition 1 (Metric Subregularity).
A set-valued map $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{q}$ is said to be metrically subregular at $(\bar{x},\bar{y})\in\mathrm{gph}(\Psi)$ with modulus $\kappa>0$ if there exists $\epsilon>0$ such that:
$\mathrm{dist}(x,\Psi^{-1}(\bar{y}))\leq\kappa\cdot\mathrm{dist}(\bar{y},\Psi(x)),\quad\forall x\in\mathbb{B}_{\epsilon}(\bar{x}).$ (6)
We consider the KKT mapping:
$T_{\text{KKT}}(w):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$, (7)
and solution set $\Omega^{*}:=\{w\mid 0\in T_{\text{KKT}}(w)\}$.
Under the above framework, we are now in position to state the following linear convergence theorem which established a Q-linear rate of distance to the solution set, and a R-linear rate of suboptimality. The proof is deferred to Appendix C.
Theorem 2.
Suppose $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{w},0)$ with modulus $c$ for any $\bar{w}\in\Omega^{*}$. Then the sequence $\{w^{(t)}\}$ generated by DS-ADMM converges Q-linearly to $\Omega^{*}$, i.e., there exist integer $T>0$ and constant $\epsilon>0$ such that for all $t>T$:
$\mathrm{dist}^{2}_{H}(w^{(t+1)},\Omega^{*})\leq\frac{1}{1+\epsilon}\cdot\mathrm{dist}^{2}_{H}(w^{(t)},\Omega^{*}),$ (8)
where
$\epsilon=\frac{\phi}{c^{2}\delta\theta}>0,\quad\phi=\min\left\{2\beta\rho,\frac{1-r}{\beta}\right\},$ (9)
and
$\delta=\max\left\{6r^{2}+\frac{2}{\beta^{2}},\,12\beta^{2}+4+(\tau\beta)^{2},\,3(\tau\beta)^{2}\right\}\quad\theta=\frac{2r^{2}\beta^{2}+1}{\beta(r+1)}+(2+\tau-r)\beta.$ (10)
Also the suboptimality converges R-linearly, which means there exists $l>0$:
$|f(u^{(t)})+g(v^{(t)})-f(u^{\infty})+g(v^{\infty})|\leq lq^{t},\quad q=\sqrt{\frac{1}{1+\epsilon}}.$ (11)
The linear convergence rate clearly depends on the algorithmic parameters $\tau$, $r$, $\beta$, and the structure of the mixing matrix $W$. In particular, a larger value of $\rho$, which reflects better network connectivity, leads to a faster convergence rate.
4.3 Sufficient Conditions for Metric Subregularity
First we state the important definition of PLQ functions:
Definition 2.
A function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is piecewise linear-quadratic (PLQ) if it is quadratic on a finite union of polyhedral regions:
$f(x)=\frac{1}{2}x^{\top}Qx+c^{\top}x+r.$ (12)
Many loss and regularization terms used in machine learning are PLQ, including the $\ell_{1}$ and $\ell_{2}$ norm, hinge loss, squared loss, and elastic net.
The following proposition gives a characterization of the metric subregularity of $T_{\mathrm{KKT}}$. Each case is justified by different results from the literature: Theorem 46 and Theorem 60 of (ADMMconv, ) support the first condition, Robinson’s continuity property (Robinson, ) establishes the second, and Lemma 4 of (PD, ) together with Theorem 60 of (ADMMconv, ) imply the third.
Proposition 4.
The KKT mapping $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{w},0)$ for any $\bar{w}\in\Omega^{*}$ under any of the following conditions: (i) each $f_{i}$ is smooth and strongly convex, and each $g_{i}$ is PLQ; (ii) all $f_{i}$ and $g_{i}$ are PLQ; (iii) all $f_{i}$ and $g_{i}$ are smooth and strongly convex.
Therefore, DS-ADMM achieves linear convergence across a wide range of practical decentralized optimization problems, including Lasso, logistic regression, SVM classification and other models frequently encountered in machine learning.
5 Numerical Experiments
We evaluate the performance of our proposed algorithm on standard decentralized composite optimization problems with smooth–nonsmooth structure, where either the objective or penalty function is nonsmooth. Comparisons are made against four representative methods: Decentralized Proximal ADMM (DPADMM, ), PG-EXTRA (PG-EXTRA, ), NIDS  (NIDS, ) and ProxMudag (ProxMudag, ). All experiments are conducted on a machine equipped with an Intel Core i7-1260P CPU and 16GB RAM.
Throughout the experiments, the number of agents is fixed at $n=30$. The communication network is modeled as a random graph with edge probability $p=0.5$, and the corresponding mixing matrix $W$ is constructed using Metropolis-Hastings weights (see Appendix A for details). Each dataset is partitioned evenly among the agents to define local loss functions. The global regularization term $g(x)$ is split uniformly as $g_{i}(x)=\frac{1}{n}g(x)$ for all agents.
All adaptive parameters (such as step sizes and penalty coefficients) are tuned to ensure the best empirical performance. Non-adaptive parameters follow the choices and guidelines suggested in the original works of the respective baseline methods. For our algorithm, we use fixed dual step sizes $(r,s)=(0.99,1)$ and a proximal coefficient $\tau=0.01$.
We measure suboptimality using $F(\bar{u}^{(t)})-F(u^{*})$, where $\bar{u}^{(t)}$ is the average of local primal variables at iteration $t$, and $u^{*}$ is a centralized optimal solution computed to high precision.
5.1 Lasso Regression
Figure 1: Performance on the Lasso regression task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
We first test the algorithms on a Lasso regression task, which consists of a quadratic loss and an $\ell_{1}$-regularization term. We use the a9a dataset from the LIBSVM repository (Libsvm, ). The objective function is given by:
$f_{i}(x)=\frac{1}{2m}\|A_{i}x-b_{i}\|^{2},\quad g_{i}(x)=\frac{\lambda}{n}\|x\|_{1},$
where $A_{i}$ and $b_{i}$ represent the local data on agent $i$, and $m$ is the total number of training samples. Here, $f_{i}$ is smooth and strongly convex, while $g_{i}$ is convex but nonsmooth. The regularization parameter is set to $\lambda=\frac{1}{m}$.
Figure 1 compares performance in terms of both iteration count and total communication rounds. Our proposed algorithm demonstrates significantly lower computational cost and reduced communication overhead compared to existing methods on this task.
5.2 SVM Classification
Next, we evaluate the algorithms on an $\ell_{2}$-regularized SVM classification problem, which involves hinge loss and an $\ell_{2}$-norm regularizer. The a1a dataset from LIBSVM is used. The objective function is:
$f_{i}(x)=\frac{1}{m}\sum_{j\in S_{i}}\max(0,1-b_{j}a_{j}^{\top}x),\quad g_{i}(x)=\frac{\lambda}{2n}\|x\|_{2}^{2},$
where $S_{i}$ is the local data index set on agent $i$, and $m$ is the total number of samples. In this setting, $f_{i}$ is convex but nonsmooth, and $g_{i}$ is smooth and strongly convex. We again use $\lambda=\frac{1}{m}$.
Note that ProxMudag is not included in this comparison because it requires the nonsmooth term to be globally coupled, which is incompatible with this separable formulation (ProxMudag, ). The results in Figure 2 show that our method achieves both faster convergence and lower communication costs.
Figure 2: Performance on the SVM classification task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Additional experimental results of graph with $p=0.2$ are provided in Appendix D. These results demonstrate that DS-ADMM consistently outperforms existing methods across both tasks and further validate the linear convergence guarantees presented in our theoretical analysis.
6 Conclusion
We proposed DS-ADMM, a fully decentralized algorithm for composite optimization based on the symmetric ADMM framework. The method integrates a novel communication structure that structures double communication per iteration without increasing overhead, while maintaining a symmetric update pattern across variable blocks. Theoretical analysis established both sublinear convergence under standard conditions, with linear rates guaranteed by metric subregularity of the KKT mapping. Our algorithm accommodates general nonsmooth regularization and is broadly applicable to various machine learning tasks. Extensive numerical experiments demonstrate that DS-ADMM outperforms existing decentralized methods in both iteration and communication efficiency.
Appendix
Appendix A Mixing Matrix Based on Metropolis-Hastings Weight
A desired mixing matrix can be constructed using Metropolis-Hastings weights, where $d_{i}=|\mathcal{N}_{i}|$ is the degree of node $i$:
$\mathcal{W}_{ij}=\begin{cases}\frac{1}{1+\max\{d_{i},d_{j}\}},&\text{if }(i,j)\in E,\\ 0,&\text{if }(i,j)\notin E\text{ and }j\neq i,\\ 1-\sum\limits_{l\in\mathcal{N}_{i}}\mathcal{W}_{il},&\text{if }j=i,\end{cases}$
Appendix B Global Form of Update
According to the update rule of Symmetric ADMM and our linear constraint formulation, a global $t$-th iteration can be written in the below equation:
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\lambda^{(t)},Au\rangle+\frac{\beta}{2}\|Au-Bv^{(t)}\|^{2}+\frac{1}{2}\|u-u^{(t)}\|_{Q}^{2},\\ \lambda^{(t+\frac{1}{2})}&=\lambda^{(t)}-r\beta(Au^{(t+1)}-Bv^{(t)})\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)+\langle\lambda^{(t+\frac{1}{2})},Bv\rangle+\frac{\beta}{2}\|Au^{(t+1)}-Bv\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ \lambda^{(t+1)}&=\lambda^{(t+\frac{1}{2})}-s\beta(Au^{(t+1)}-Bv^{(t+1)}).\end{aligned}\right$. (13)
Then we decompose the Lagrange multiplier $\lambda$:
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle w_{1}^{(t)},\tilde{W}u\rangle-\langle w_{2}^{(t)},u\rangle+\frac{\beta}{2}\|\tilde{W}u-v^{(t)}\|^{2}+\frac{\beta}{2}\|u-\tilde{W}v^{(t)}\|^{2}\\ &+\frac{1}{2}\|u-u^{(t)}\|_{Q}^{2},\\ w_{1}^{(t+\frac{1}{2})}&=w_{1}^{(t)}-r\beta(\tilde{W}u^{(t+1)}-v^{(t)}),\quad w_{2}^{(t+\frac{1}{2})}=w_{2}^{(t)}-r\beta(u^{(t+1)}-\tilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)+\langle w_{1}^{(t+\frac{1}{2})},v\rangle+\langle w_{2}^{(t+\frac{1}{2})},\tilde{W}v\rangle+\frac{\beta}{2}\|\tilde{W}u^{(t+1)}-v\|^{2}\\ &+\frac{\beta}{2}\|u^{(t+1)}-\tilde{W}v\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ w_{1}^{(t+1)}&=w_{1}^{(t+\frac{1}{2})}-s\beta(\tilde{W}u^{(t+1)}-v^{(t+1)}),\quad w_{2}^{(t+1)}=w_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\tilde{W}v^{(t+1)}).\end{aligned}\right$. (14)
which can be transformed into the following version ready for decentralized update:
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\tilde{W}w_{1}^{(t)}+w_{2}^{(t)}+\beta(2\tilde{W}v^{(t)}+(1+\tau)u^{(t)}-\tilde{W}^{\top}\tilde{W}u^{(t)}),u\rangle\\ &+\beta(1+\frac{\tau}{2})\|u\|^{2}\\ w_{1}^{(t+\frac{1}{2})}&=w_{1}^{(t)}-r\beta(\tilde{W}u^{(t+1)}-v^{(t)}),\quad w_{2}^{(t+\frac{1}{2})}=w_{2}^{(t)}-r\beta(u^{(t+1)}-\tilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)-\langle\beta(2\tilde{W}u^{(t+1)}+(1+\tau)v^{(t)}-\tilde{W}^{\top}\tilde{W}v^{(t)})-(w_{1}^{(t+\frac{1}{2})}+\\ &\tilde{W}w_{2}^{(t+\frac{1}{2})}),v\rangle+\beta(1+\frac{\tau}{2})\|v\|^{2}\\ w_{1}^{(t+1)}&=w_{1}^{(t+\frac{1}{2})}-s\beta(\tilde{W}u^{(t+1)}-v^{(t+1)}),\quad w_{2}^{(t+1)}=w_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\tilde{W}v^{(t+1)}).\end{aligned}\right$. (15)
Appendix C Proof of Theorem 2
The proof is based on the notations of global form (5). First, we denote vectors
$z=\begin{pmatrix}u\\ v\end{pmatrix},\quad w=\begin{pmatrix}u\\ v\\ \lambda\end{pmatrix},\quad F(w)=\begin{pmatrix}-A^{\top}\lambda\\ B^{\top}\lambda\\ Au-Bv\end{pmatrix}$ (16)
$\widetilde{u}^{(t)}=u^{(t+1)},\quad\widetilde{v}^{(t)}=v^{(t+1)},\quad\widetilde{\lambda}^{(t)}=\lambda^{(t)}-\beta\left(Au^{(t+1)}-Bv^{(t+1)}\right),$ (17)
$\widetilde{z}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\end{pmatrix},\quad\widetilde{w}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\\ \widetilde{\lambda}^{(t)},\end{pmatrix},\quad h(z)=f(x)+g(y)$ (18)
and matrices
$S=\begin{pmatrix}Q&&\\ &Q+\beta B^{\top}B&-rB^{\top}\\ &B&\frac{1}{\beta}I\end{pmatrix},M=\begin{pmatrix}I&&\\ &I&\\ &-\beta B&(1+r)I\end{pmatrix}$, (19)
$H=\begin{pmatrix}Q&&\\ &Q+\frac{1}{r+1}\beta B^{\top}B&-\frac{r}{r+1}B^{\top}\\ &-\frac{r}{r+1}B&\frac{1}{\beta(r+1)}I\end{pmatrix},G=\begin{pmatrix}Q&&\\ &Q&\\ &&\frac{1-r}{\beta}I\end{pmatrix}$, (20)
It is easy to verify the following properties:
$w^{(t+1)}=w^{(t)}-M(w^{(t)}-\widetilde{w}^{(t)}).$ (21)
$G=S+S^{\top}-M^{\top}S,\quad H=SM^{-1}.$ (22)
The following lemma characterizes the eigenvalue property of $H$ and $G$:
Lemma 1.
$\lambda_{\max}(H)\leq\theta,\quad\lambda_{\min}(G)\geq 2(1-r)\rho$ (23)
Proof.
$H=\begin{pmatrix}Q&&\\ &Q+(1-r)\beta B^{\top}B+\frac{r^{2}}{r+1}\beta B^{\top}B&-\frac{r}{r+1}B^{\top}\\ &-\frac{r}{r+1}B&\frac{1}{\beta(r+1)}I\end{pmatrix}$ (24)
so it is easy to verify that
$\lambda_{\max}(H)\leq\lambda_{\max}(Q+(1-r)\beta B^{\top}B)+\lambda_{\max}(\frac{1}{\beta(r+1)}\begin{pmatrix}r\beta B^{\top}\\ -I\end{pmatrix}\begin{pmatrix}r\beta B^{\top}\\ -I\end{pmatrix}^{\top})$ (25)
while
$\lambda_{\max}(Q+(1-r)\beta B^{\top}B)=\lambda_{\max}(\beta(1+\tau)I-\beta\widetilde{W}^{\top}\widetilde{W}+(1-r)\beta(I+\widetilde{W}^{\top}\widetilde{W}))\leq(2+\tau-r)\beta,$ (26)
$\lambda_{\max}(\frac{1}{\beta(r+1)}\begin{pmatrix}r\beta B^{\top}\\ -I\end{pmatrix}\begin{pmatrix}r\beta B^{\top}\\ -I\end{pmatrix}^{\top})=\frac{1}{\beta(r+1)}\lambda_{\max}((r\beta)^{2}(I+\widetilde{W}^{\top}\widetilde{W})+I)=\frac{2r^{2}\beta^{2}+1}{\beta(r+1)},$ (27)
combining above, we get $\lambda_{\max}(H)\leq\theta$. Also, we have
$\lambda_{\min}(G)=\min\{\lambda_{\min}(Q),\frac{1-r}{\beta}\}\geq\min\{2\beta\rho,\frac{1-r}{\beta}\}=\phi.$ (28)
This equivalent form of solution set $\Omega^{*}$ follows from [14] by the definition of subdifferentials:
$\Omega^{*}=\bigcap_{w\in\mathbb{R}^{4nd}}\left\{\widehat{w}\,\middle|\,h(z)-h(\widehat{z})+\langle w-\widehat{w},F(w)\rangle\geq 0\right\}.$ (29)
Then we give the following important lemma, whose proof follows directly from the proof of Theorem 2 and Theorem 3 in [3] and the above form of solution set:
Lemma 2.
$\left\|w^{(t+1)}-w^{*}\right\|_{H}^{2}\leq\left\|w^{(t)}-w^{*}\right\|_{H}^{2}-\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2},\quad\forall w^{*}\in\Omega^{*},$ (30)
Let us revise the KKT mapping
$T_{\text{KKT}}(w):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$. (31)
The following lemma bounds the distance through $G$-norm.
Lemma 3.
The sequences $\{w^{(t)}\}$ and $\{\widetilde{w}^{(t)}\}$ satisfy
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w^{(t+1)}))\leq\frac{\delta}{\theta}\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2},$
Proof.
By Proposition 2.1 in [14] and our notations, we derive the following inequality characterizing the subproblems:
$f(u)-f(\widetilde{u}^{(t)})+\left\langle u-\widetilde{u}^{(t)},-A^{\top}\widetilde{\lambda}^{(t)}+Q(\widetilde{u}^{(t)}-u^{(t)})\right\rangle\geq 0$ (32)
$g(v)-g(\widetilde{v}^{(t)})+\left\langle v-\widetilde{v}^{(t)},-B^{\top}\widetilde{\lambda}^{(t)}-rB^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t)})+(Q+\beta B^{\top}B)(\widetilde{v}^{(t)}-v^{(t)})\right\rangle\geq 0$ (33)
It is obvious that
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w))=\mathrm{dist}^{2}(0,T_{1}(w))+\mathrm{dist}^{2}(0,T_{2}(w))+\mathrm{dist}^{2}(0,T_{3}(w))$ (34)
where $T_{1}(w)=\partial f(u)-A^{\top}\lambda$, $T_{2}(w)=\partial g(v)+B^{\top}\lambda$, $T_{3}(w)=Au-Bv$.
From the above two inequalities, we have the following:
$\mathrm{dist}(0,T_{1}(w^{(t+1)}))$ $\leq\left\|A^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|$ (35)
$=\left\|A^{\top}\left[r(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-\beta B(v^{(t)}-\widetilde{v}^{(t)})\right]-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|,$
where the second equality uses the update rule
$\lambda^{(t+1)}$ $=\lambda^{(t+\frac{1}{2})}-\beta(Au^{(t+1)}-Bv^{(t+1)})$ (36)
$=\lambda^{(t+\frac{1}{2})}-\beta(Au^{(t+1)}-Bv^{(t+1)})+\beta B(v^{(t)}-v^{(t+1)})$
$=\lambda^{k}-(r+1)(\lambda^{k}-\widetilde{\lambda}^{k})+\beta B(v^{(t)}-v^{(t+1)})$
similarly, we have:
$\mathrm{dist}(0,T_{2}(w^{(t+1)}))$ $\leq\left\|B^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-(Q+\beta B^{\top}B)(\widetilde{v}^{(t)}-v^{(t)})+rB^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t)})\right\|$ (37)
$=\left\|Q(\widetilde{v}^{(t)}-v^{(t)})\right\|,$
and finally
$\mathrm{dist}(0,T_{3}(w^{(t+1)}))$ $=\left\|\frac{1}{\beta}(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-B(v^{(t)}-\widetilde{v}^{(t)})\right\|.$ (38)
Substituting  (35) (37) (38) into (34) while using the definition of $A$, $B$, $Q$, we get
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w^{(t+1)}))$ $\leq\left\|A^{\top}\left[r(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-\beta B(v^{(t)}-\widetilde{v}^{(t)})\right]-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|^{2}$ (39)
$+\left\|Q(\widetilde{v}^{(t)}-v^{(t)})\right\|^{2}+\left\|\frac{1}{\beta}(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-B(v^{(t)}-\widetilde{v}^{(t)})\right\|^{2}$
$\leq 3r^{2}\left\|A^{\top}(\lambda^{(t)}-\widetilde{\lambda}^{(t)})\right\|^{2}+\frac{2}{\beta^{2}}\left\|\lambda^{(t)}-\widetilde{\lambda}^{(t)}\right\|^{2}$
$+3\beta^{2}\left\|A^{\top}B(v^{(t)}-\widetilde{v}^{(t)})\right\|^{2}$
$+2\left\|B(v^{(t)}-\widetilde{v}^{(t)})\right\|^{2}+\left\|Q(v^{(t)}-\widetilde{v}^{(t)})\right\|^{2}+3\left\|Q(\widetilde{u}^{(t)}-u^{(t)})\right\|^{2}$
[…]
$+3(\tau\beta)^{2}\left\|u^{(t)}-\widetilde{u}^{(t)}\right\|^{2}$
$\leq\delta\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|^{2}\ \leq\frac{\delta}{\phi}\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2}.$
Then we are able to prove the first part of Theorem 2:
Proof.
Because $\Omega^{*}$ is a closed convex set, there exists a $w^{*}_{t}\in\Omega^{*}$ satisfying
$\operatorname{dist}_{H}(w^{(t)},\Omega^{*})=\left\|w^{(t)}-w^{*}_{t}\right\|_{H}.$ (40)
Then, by the metric subregularity of the KKT mapping and the global convergence result, there exists $T>0$, for any $t>T$ we have
$\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}\geq$ $\sqrt{\frac{\phi}{\delta}}\mathrm{dist}(0,T_{\text{KKT}}(w^{(t+1)}))$ (41)
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta}}\operatorname{dist}(w^{(t+1)},\Omega^{*})$
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta\theta}}\operatorname{dist}_{H}(w^{(t+1)},\Omega^{*}).$
So, we will have from the above inequality that
$(1+\epsilon)\operatorname{dist}_{H}^{2}(w^{(t+1)},\Omega^{*})\leq$ $\left\|w^{(t+1)}-w^{*}_{t}\right\|_{H}^{2}+\epsilon\operatorname{dist}_{H}^{2}(w^{(t+1)},\Omega^{*})$ (42)
$\leq$ $\left\|w^{(t+1)}-w^{*}_{t}\right\|_{H}^{2}+\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2}$
$\leq$ $\left\|w^{(t)}-w^{*}_{t}\right\|_{H}^{2}=\operatorname{dist}_{H}^{2}(w^{(t)},\Omega^{*})$
and the main result is proved.
To prove the R-linear rate of suboptimality, first, from the same approach of proving Corollary 2.1 in [2] we are able to prove the R-linear rate of $\left\|w^{(t)}-w^{\infty}\right\|$. And a simple corollary from Lemma 2 is that $\left\|w^{(t)}-\tilde{w}^{(t)}\right\|$ and $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ also converges R-linearly.
Then we substitute $u^{\infty}$ and $v^{\infty}$ into (32) and  (33):
$f(u^{\infty})+g(v^{\infty})$ $\geq f(\widetilde{u}^{(t)})+g(\widetilde{v}^{(t)})-\left\langle u^{\infty}-\widetilde{u}^{(t)},-A^{\top}\widetilde{\lambda}^{(t)}+Q(\widetilde{u}^{(t)}-u^{(t)})\right\rangle$ (43)
$-\left\langle v^{\infty}-\widetilde{v}^{(t)},-B^{\top}\widetilde{\lambda}^{(t)}-rB^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t)})+(Q+\beta B^{\top}B)(\widetilde{v}^{(t)}-v^{(t)})\right\rangle$
and from the definition of solution set
$f(u^{(t+1)})+g(v^{(t+1)})\geq f(u^{\infty})+g(v^{\infty})+\langle\lambda^{\infty},Au^{(t+1)}-Bv^{(t+1)}\rangle.$ (44)
the above inequalities and the R-linear convergence of $\left\|w^{(t)}-w^{\infty}\right\|$, $\left\|w^{(t)}-\tilde{w}^{(t)}\right\|$, $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ lead to the R-linear convergence of suboptimality, which finishes the proof.
Appendix D Additional Experimental Results
In this section, we report experimental results on a 30-agent sparsely connected random graph with edge probability $p=0.2$. As expected, all algorithms perform worse on this sparsely connected topology, consistent with the theoretical results. Nevertheless, our proposed DS-ADMM still demonstrates the best overall performance.
D.1 Lasso Regression
Figure 3: Performance on the Lasso regression task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
D.2 SVM Classification
Figure 4: Performance on the SVM classification task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.